OpenStudy (anonymous):

chemistry help Convert 1.03 g/mL to the unit g/m^3.

OpenStudy (anonymous):

first you go to liters 1 liter = 1000ml so 1.03g/ml = 1030g/L then 1 L = 1 dm^3 so you get 1030 g/dm^3 1 m ^3 = 1000dm^3 so you get 1,030,000 g/m^3

OpenStudy (anonymous):

@aerospacenerd

OpenStudy (anonymous):

the answers are not like that

OpenStudy (anonymous):

look

OpenStudy (anonymous):

1,030,000 = 1.03*10^6

OpenStudy (anonymous):

ok